OpenStudy (aravindg):

the value of the expression 1-(sin^2y/[1+cos y])+([1+cos y]/siny)-(sin y/[1-cos y])=????

myininaya (myininaya):

can you use the equation editor

OpenStudy (aravindg):

kkk

myininaya (myininaya):

wait maybe i can make it out

OpenStudy (aravindg):

...................

myininaya (myininaya):

cos(y)

OpenStudy (aravindg):

\[1-(\sin ^{2}y \div1+\cos y)+(1+\cos y \div \sin y)-(\sin y \div1-\cos y)\]

OpenStudy (aravindg):

howwww??????????????????

myininaya (myininaya):

give me just a second

OpenStudy (aravindg):

is tht much step necessary????

myininaya (myininaya):

i don't know i was playing with

OpenStudy (aravindg):

can u giv short steps???

OpenStudy (aravindg):

i need quicker way

myininaya (myininaya):

just combine fractions earlier instead of multiplying by conjugates

OpenStudy (anonymous):

I got cos y

OpenStudy (aravindg):

show

OpenStudy (anonymous):

that pdf posted above is the complete solution, no shortcuts. Just remember that sin^2 y + cos^2 y = 1 &&& (1+cosy)(1-cosy) = (1-cos^2 y)

myininaya (myininaya):

\[1-\frac{\sin^2(y)}{1+\cos(y)} \cdot \frac{1-\cos(y)}{1-\cos(y)}+\frac{(1+\cos(y))(1-\cos(y))-\sin(y)\sin(y)}{\sin(y)(1-\cos(y))}\] \[1-\frac{\sin^2(y)(1-\cos(y))}{1-\cos^2(y)}+\frac{1-\cos^2(y)-\sin^2(y)}{\sin(y)(1-\cos(y))}\] \[1-\frac{\sin^2(y)(1-\cos(y))}{\sin^2(y)}+\frac{\sin^2(y)-\sin^2(y)}{\sin(y)(1-\cos(y))}\] \[1-(1-\cos(y))+\frac{0}{\sin(y)(1-\cos(y))}\] \[1-1+\cos(y)+0=\cos(y)\]

OpenStudy (anonymous):

Don't forget that y cannot be 180 degrees or 0 degrees!

myininaya (myininaya):

\[\sin(y) \neq0 or 1-\cos(y) \neq 0\] \[y \neq n \pi, n \in \mathbb{Z} \]

OpenStudy (anonymous):

^ Or you can use overly complicated notation :D Great job ininaya :D

myininaya (myininaya):

lol

OpenStudy (anonymous):

Is there a way I can delete my medal? I don't deserve it :D

myininaya (myininaya):

well i was thinking about saying when that bottom is not zero but you said it for me when i didn't feel like it so you deserve it because what you said is true

OpenStudy (anonymous):

You are awesome. PERIOD (not monthly discharge).